A 2.0 uf capacitor in a circuit in series with a resistance of 1.0 MO is charged with a 6.0 Volts vatery. How long would it take to charge the capacitor to 3/4 of its maximum voltage?
Answer #1
In a capacitor,

V = Vo* ( 1- e^(-t/RC))
or, 3Vo/4 =Vo* ( 1- e^(-t/RC))
or, 3/4 = 1- e^(-t/RC)
or, e^(t/RC) =4
t = RC ln 4 = 2.7725 seconds